# Metrics on completely positive maps via noncommutative geometry

### v1

@@ 6.2 Group $C^{*}$-algebras and length functions
Let $G$ be a discrete group, $l\colon G\to[0,\infty)$ be a length function, and $\sigma\in Z^{2}(G,\mathbb{T})$. Suppose $\phi\in P(G)$. Then $M_{\phi}$ extends to a bounded multiplier $T^{\phi}$ of the twisted crossed product $\ell^{\infty}(G)\rtimes_{\mathrm{red},\sigma}G$, and we have
Proof.
The fact that $M_{\phi}$ extends to a completely positive map on $\ell^{\infty}(G)\rtimes_{\mathrm{red},\sigma}G$ satisfying
$T^{\phi}(\psi\lambda^{\sigma}_{g})=\phi(g)\psi\lambda^{\sigma}_{g}\quad\text{ for all }\psi\in\ell^{\infty}(G)\text{ and }g\in G,$ (6.5)
Hence, for each $f\in C_{c}(G,\sigma)$ we get
$=\|T^{\phi}([\partial_{l},\lambda^{\sigma}(f)])\|$
$\leq\|T^{\phi}\|L_{\partial_{l}}(\lambda^{\sigma}(f))$
$=\|M_{\phi}\|\,L_{\partial_{l}}(\lambda^{\sigma}(f)).$
Lemma 6.6.
Let $G$ be a discrete group, $\sigma\in Z^{2}(G,\mathbb{T})$, and $\phi\in P(G)$. Then the completely positive map $M_{\phi}\colon C^{*}_{r}(G,\sigma)\to C^{*}_{r}(G,\sigma)$ is $(\tau_{\sigma},\tau_{\sigma})$-adjointable, and $M_{\phi}^{\sharp}=M_{\phi^{\circ}}$ where $\phi^{\circ}\in P(G)$ is given by
@@
That $\phi^{\circ}\in P(G)$ follows by [BC16, p. 7]. Hence, $M_{\phi^{\circ}}\colon C^{*}_{r}(G,\sigma)\to C^{*}_{r}(G,\sigma)$ is completely positive. For $f_{1},f_{2}\in C^{*}_{r}(G,\sigma)$ we get that
@@
This implies that $M_{\phi}$ is $(\tau_{\sigma},\tau_{\sigma})$-adjointable, with $M_{\phi}^{\sharp}=M_{\phi^{\circ}}$. ∎
@@
Let $G$ be a countable discrete amenable group, $l\colon G\to[0,\infty)$ be a proper length function, and $\sigma\in Z^{2}(G,\mathbb{T})$. Then the metric $\Delta_{\tau_{\sigma},L_{\partial_{l}\times\partial_{l}}}^{\rm min}$ induced from the exterior Kasparov product of $(C^{*}_{r}(G,\sigma),\ell^{2}(G),\partial_{l})$ with $(C^{*}_{r}(G,\sigma)^{\rm op},\ell^{2}(G),\partial_{l})$ restricted to $P_{1}(G)$ satisfies chaining. That is
@@
By [Béd95, Corollary 3.8] we have that if $G$ is amenable then $\tau_{\sigma}$ is an amenable tracial state, hence the metric $\Delta_{\tau_{\sigma},L_{\partial_{l}\times\partial_{l}}}^{\rm min}$ is well-defined. To apply Theorem 5.10 and Remark 5.11 we only need to verify that
@@
for every $\phi\in P_{1}(G)$. As in the proof of Lemma 6.5 we compute the following for $x\in C^{*}_{r}(G,\sigma)\otimes_{\rm min}C^{*}_{r}(G,\sigma)^{\rm op}$ and $\phi\in P_{1}(G)$:
$\left[\begin{pmatrix}0&\partial_{l}\otimes 1+i\otimes\partial_{l}\\ \partial_{l}\otimes 1+i\otimes\partial_{l}&0\end{pmatrix},\begin{pmatrix}(M_{\phi}\otimes\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}})(x)&0\\ 0&(M_{\phi}\otimes\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}})(x)\end{pmatrix}\right$]
@@
Computing the operator norm of this proves that $L_{\partial_{l}\times\partial_{l}}\circ(M_{\phi}\otimes\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}}))\leq L_{\partial_{l}\times\partial_{l}}$, since
@@
$\Sigma^{\rm op}((M_{\phi}\otimes\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}})(x))=\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}}\otimes M_{\phi}^{\rm op}(\Sigma^{\rm op}(x)).$
We furthermore have by Lemma 6.6 that $M_{\phi}^{\sharp}=M_{\phi^{\circ}}$, Hence it suffices to show that
$L_{\partial_{l}\times\partial_{l}}\circ(\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}})\otimes M_{\phi}^{\rm op})\leq L_{\partial_{l}\times\partial_{l}}$
@@
Combining Theorem 6.2 and Theorem 6.9 gives many examples of stable metrics that satisfy chaining. We record the following example explicitly.
@@
Let $G$ be a countable discrete amenable group, $\sigma\in Z^{2}(G,\mathbb{T})$, and $l\colon G\to[0,\infty)$ be a proper length function. As $G$ is amenable, $C^{*}_{r}(G,\sigma)$ is a nuclear $C^{*}$-algebra. This yields spectral triples $(C^{*}_{r}(G),\ell^{2}(G),\partial_{l})$ and $(C^{*}_{r}(G,\sigma)^{\rm op},\ell^{2}(G),\partial_{l})$ as we have seen above. We will obtain a stable sequence of metrics $\{\Delta_{\operatorname{Tr}_{n}\otimes\tau,L_{n}}^{\rm min}\}_{n\in\mathbb{N}}$ by choosing a spectral triple $(M_{n}(\mathbb{C}),H_{n},\partial_{n})$ for each $n\in\mathbb{N}$ by Theorem 6.2. Such spectral triples have been studied in [DM21], where they show how to recover the matricial Wasserstein-1 distance from [Che+17] from the point of view of noncommutative geometry. Namely, for each $n\in\mathbb{N}$ pick $N\in\mathbb{N}$, set $H_{n}=\mathbb{C}^{n}\otimes\mathbb{C}^{N}$ and
$\partial_{n}=\sum_{i=1}^{N}L_{i}\otimes e_{i,i}$
where $L_{i}\in M_{n}(\mathbb{C})$ is a self-adjoint matrix for each $i=1,\ldots,N$. We then have a spectral triple $(M_{n}(\mathbb{C}),H_{n},\partial_{n})$ for each $n\in\mathbb{N}$. Recall that for any $\varphi\in S(M_{n}(\mathbb{C}))$, there is $\rho\in M_{n}(\mathbb{C})$ such that $\varphi=\operatorname{Tr}_{n}(\rho\,\cdot)$. D’Andrea and Martinetti show in [DM21, Proposition 5] that the associated Monge-Kantorovič metric can be realized as
$\operatorname{mk}_{\partial_{n}}(\varphi_{1},\varphi_{2})=\inf\{\operatorname{Tr}_{n}(\sqrt{u^{*}u})\mid u_{i}\in M_{n}(\mathbb{C}),\sum_{i=1}^{N}[L_{i},u_{i}]=\rho_{1}-\rho_{2}\}$
where $\varphi_{j}=\operatorname{Tr}_{n}(\rho_{j}\cdot)$. These spectral triples are therefore related to the trace norm. For each $n$, make such a choice of spectral triple $(M_{n}(\mathbb{C}),H_{n},\partial_{n})$. We then obtain a sequence of seminorms $L_{n}=L_{(\partial_{n}\times\partial_{n})\times(\partial_{l}\times\partial_{l})}\circ\Sigma_{[23]}$ by taking the appropriate Kasparov products as in Theorem 6.2, which yield a stable sequence of metrics $\{\Delta_{\operatorname{Tr}_{n}\otimes\tau,L_{n}}^{\rm min}\}_{n\in\mathbb{N}}$, where each $\Delta_{\operatorname{Tr}_{n}\otimes\tau,L_{n}}^{\rm min}$ is a metric on $\operatorname{TC}_{\operatorname{Tr}_{n}\otimes\tau}(M_{n}(\mathbb{C})\otimes C^{*}_{r}(G,\sigma),M_{n}(\mathbb{C})\otimes C^{*}_{r}(G,\sigma))$. By Theorem 6.9, the metric $\Delta_{\operatorname{Tr}_{1}\otimes\tau_{\sigma},L_{1}}^{\rm min}=\Delta_{\tau_{\sigma},L_{\partial_{l}\times\partial_{l}}}^{\rm min}$ satisfies chaining with respect to the unital completely positive maps arising as multipliers $M_{\phi}$, $\phi\in P_{1}(G)$.

### v2

@@ 7.2. Amenable groups and twisted group $C^{*}$-algebras
Let $G$ be a discrete group, $\sigma\in Z^{2}(G,\mathbb{T})$, and $\phi\in P(G)$. Then the completely positive map $M_{\phi}\colon C^{*}_{r}(G,\sigma)\to C^{*}_{r}(G,\sigma)$ is $(\tau_{\sigma},\tau_{\sigma})$-adjointable, and $M_{\phi}^{\sharp}=M_{\phi^{\circ}}$ where the function $\phi^{\circ}\in P(G)$ is given by
@@
That $\phi^{\circ}\in P(G)$ follows by [BC16, p. 7]. Hence, $M_{\phi^{\circ}}\colon C^{*}_{r}(G,\sigma)\to C^{*}_{r}(G,\sigma)$ is completely positive. For $f_{1},f_{2}\in C_{c}(G,\sigma)$, we get that
@@
By continuity, this implies that $M_{\phi}$ is $(\tau_{\sigma},\tau_{\sigma})$-adjointable, with $M_{\phi}^{\sharp}=M_{\phi^{\circ}}$. ∎
@@
Let $G$ be a countable discrete amenable group, $l\colon G\to[0,\infty)$ be a proper length function, and $\sigma\in Z^{2}(G,\mathbb{T})$. Then the metric $\Delta_{\tau_{\sigma},L_{\partial_{l}\times\partial_{l}}}^{\rm min}$ induced from the external Kasparov product of $(C^{*}_{r}(G,\sigma),\ell^{2}(G),\partial_{l})$ with $(C^{*}_{r}(G,\sigma)^{\rm op},\ell^{2}(G),\partial_{l})$ restricted to $P_{1}(G)$ satisfies chaining. That is, we have
@@
By [B(Font) ̵ ̵ ̵ ̵ ̵ ̵ ̵ ̵ ̵ ̵ ̵ ̵ ̵ ̵LaTeX ̵Font ̵Info: ̵ ̵ ̵ ̵éd95, Corollary 3.8] we have that if $G$ is amenable then $\tau_{\sigma}$ is an amenable tracial state, hence the metric $\Delta_{\tau_{\sigma},L_{\partial_{l}\times\partial_{l}}}^{\rm min}$ is well-defined. To apply theorem 6.11 and remark 6.12 we only need to verify that
@@
for every $\phi\in P_{1}(G)$. Let $x\in C^{*}_{r}(G,\sigma)\otimes_{\rm min}C^{*}_{r}(G,\sigma)^{\rm op}$ and $\phi\in P_{1}(G)$. As in the proof of lemma 2.12, we compute that
$\left[\begin{pmatrix}0&\partial_{l}\otimes 1+i\otimes\partial_{l}\\ \partial_{l}\otimes 1-i\otimes\partial_{l}&0\end{pmatrix},\begin{pmatrix}(M_{\phi}\otimes\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}})(x)&0\\ 0&(M_{\phi}\otimes\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}})(x)\end{pmatrix}\right$]
@@
Computing the operator norm of this proves that $L_{\partial_{l}\times\partial_{l}}\circ(M_{\phi}\otimes\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}})\leq L_{\partial_{l}\times\partial_{l}}$, since
@@
$\Sigma^{\rm op}((M_{\phi}\otimes\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}})(x))=(\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}}\otimes M_{\phi}^{\rm op})(\Sigma^{\rm op}(x)).$
We furthermore have by lemma 7.4 that $M_{\phi}^{\sharp}=M_{\phi^{\circ}}$. Hence it suffices to show that
$L_{\partial_{l}\times\partial_{l}}\circ(\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}}\otimes M_{\phi}^{\rm op})\leq L_{\partial_{l}\times\partial_{l}}$
@@
Combining theorem 7.2 and theorem 7.7 gives many examples of stable sequences of extended metrics that satisfy chaining. We record the following example explicitly.
@@
Let $G$ be a countable amenable discrete group, $\sigma\in Z^{2}(G,\mathbb{T})$, and $l\colon G\to[0,\infty)$ be a proper length function. This yields spectral triples $(C^{*}_{r}(G),\ell^{2}(G),\partial_{l})$ and $(C^{*}_{r}(G,\sigma)^{\rm op},\ell^{2}(G),\partial_{l})$ as we have seen above. As $G$ is amenable, $C^{*}_{r}(G,\sigma)$ is a nuclear $C^{*}$-algebra. We will obtain a stable sequence of metrics $\{\Delta_{\operatorname{Tr}_{n}\otimes\tau,L_{n}}^{\rm min}\}_{n\in\mathbb{N}}$ by choosing a spectral triple $(M_{n}(\mathbb{C}),H_{n},\partial_{n})$ for each $n\in\mathbb{N}$ by theorem 7.2. Such spectral triples have been studied in [DM21], where they show how to recover the matricial Wasserstein-1 distance from [Che+17] from the point of view of noncommutative geometry. Namely, pick $N\in\mathbb{N}$ and, for each $n\in\mathbb{N}$, set $H_{n}=\mathbb{C}^{n}\otimes\mathbb{C}^{N}$ and
$\partial_{n}=\sum_{i=1}^{N}R_{i}\otimes e_{i,i}^{(N)}$
where $R_{i}\in M_{n}(\mathbb{C})$ is a self-adjoint matrix for each $i=1,\ldots,N$. Letting $M_{n}(\mathbb{C})$ act on the first factor of $H_{n}$, we then have a spectral triple $(M_{n}(\mathbb{C}),H_{n},\partial_{n})$ for each $n\in\mathbb{N}$. Recall that for any $\varphi\in S(M_{n}(\mathbb{C}))$, there is $\rho\in M_{n}(\mathbb{C})$ such that $\varphi=\operatorname{Tr}_{n}(\rho\,\cdot)$. D’Andrea and Martinetti show in [DM21, Proposition 5] that the associated Monge-Kantorovič metric can be realized as
$\operatorname{mk}_{\partial_{n}}(\varphi_{1},\varphi_{2})=\inf\big\{\operatorname{Tr}_{n}(\sqrt{u^{*}u})\mid u_{i}\in M_{n}(\mathbb{C}),\sum_{i=1}^{N}[R_{i},u_{i}]=\rho_{1}-\rho_{2}\big\}$
where $\varphi_{j}=\operatorname{Tr}_{n}(\rho_{j}\cdot)$ and $u=(u_{1},\ldots,u_{N})$. These spectral triples are therefore related to the trace norm. For each $n$, make such a choice of spectral triple $(M_{n}(\mathbb{C}),H_{n},\partial_{n})$. We then obtain a sequence of seminorms $L_{n}=L_{(\partial_{n}\times\partial_{n})\times(\partial_{l}\times\partial_{l})}\circ\Sigma_{[23]}$ by taking the appropriate Kasparov products as in theorem 7.2, which yield a stable sequence of metrics $\{\Delta_{\operatorname{Tr}_{n}\otimes\tau,L_{n}}^{\rm min}\}_{n\in\mathbb{N}}$, where each $\Delta_{\operatorname{Tr}_{n}\otimes\tau,L_{n}}^{\rm min}$ is a metric on $\operatorname{TC}_{\operatorname{Tr}_{n}\otimes\tau}(M_{n}(\mathbb{C})\otimes C^{*}_{r}(G,\sigma),M_{n}(\mathbb{C})\otimes C^{*}_{r}(G,\sigma))$. By theorem 7.7, the metric $\Delta_{\operatorname{Tr}_{1}\otimes\tau_{\sigma},L_{1}}^{\rm min}=\Delta_{\tau_{\sigma},L_{\partial_{l}\times\partial_{l}}}^{\rm min}$ satisfies chaining with respect to the unital completely positive maps arising as multipliers $M_{\phi}$ for $\phi\in P_{1}(G)$.